OpenStudy (anonymous):

Find x^2-2x-4=0 using the quadratic formula. I keep getting the wrong answer. I know what the right answer is.

OpenStudy (anonymous):

\[x = 1 \pm \sqrt{5} \]

OpenStudy (anonymous):

\[x \approx 3.2360679775, ~-1.2360679775\]

OpenStudy (anonymous):

How did you get that answer? I keep getting \[\pm2\sqrt{5}\]

OpenStudy (anonymous):

Can you show your work? @Rachella

OpenStudy (anonymous):

Step One - Find coefficients a, b and c. \[a = 1, ~b= -2,~c=-4\] Step Two - Plug in the values for a, b and c into the quadratic formula. \[x1,2 = \frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a}\] \[x1,2 = \frac{ -2\pm \sqrt{(-2^{2}-4\times1\times(-4)} }{ 2\times1}\] Step Three - Simplify expression under the square root. \[x1,2 = \frac{ 2\pm \sqrt{20} }{ 2 }\] Step Four - Solve for x. \[x1 = \frac{ 2+ \sqrt{20} }{ 2 }~= 1 + \sqrt{5}\] \[x2 = \frac{ 2- \sqrt{20} }{ 2 } = 1 - \sqrt{5}\]

OpenStudy (anonymous):

\[Equation~To~Solve: x ^{2} - 2x - 4 = 0 \] \[The~Solutions~Are: x1 = 1 + \sqrt{5}~ and ~x2= 1 - \sqrt{5}\]

OpenStudy (anonymous):

How does the \[\sqrt{20}\] equal \[\sqrt{5}\]

OpenStudy (anonymous):

The square root of 20 would be 4.472 which rounded would be 5.

OpenStudy (anonymous):

When I simplified \[\sqrt{20}\] , I got \[2\sqrt{5}\]

OpenStudy (anonymous):

I agree with @Rachella

OpenStudy (anonymous):

I looked in the back of the book and Rachella's answer is correct but I am not sure how she got there.... @Rachella @iKayla

OpenStudy (anonymous):

I explained step by step how I got that answer.

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

How did you go from \[\frac{ 2+\sqrt{20} }{ 2 }\] to \[1 + \sqrt{5}\]

OpenStudy (anonymous):

\[\sqrt{20} = \sqrt{4 \times 5} = \color{green}{2 \sqrt{5}}\]

OpenStudy (anonymous):

Take out \(2\) common from Numerator and cancel it with \(2\) in denominator.

OpenStudy (anonymous):

\[\frac{2 + 2 \sqrt{5}}{2} \implies \frac{\cancel{2}(1+\sqrt{5})}{\cancel{2}} = 1 + \sqrt{5}\]

OpenStudy (anonymous):

Thanks so much! It makes more sense now! @waterineyes :)

OpenStudy (anonymous):

you are welcome dear.. :)

OpenStudy (anonymous):

You basically factored out the two right? @waterineyes

OpenStudy (anonymous):

Taking something common means yes factoring out.. I have factored out \(2\) there.

OpenStudy (anonymous):

Ok! Thanks again! :)